# MedMax: Mixed-Modal Instruction Tuning for Training Biomedical Assistants

### v1

@@ Abstract
Recent advancements in mixed-modal generative models have enabled flexible integration of information across image-text content. These models have opened new avenues for developing unified biomedical assistants capable of analyzing biomedical images, answering complex questions about them, and predicting the impact of medical procedures on a patient’s health. However, existing resources face challenges such as limited data availability, narrow domain coverage, and restricted sources (e.g., medical papers). To address these gaps, we present MedMax, the first large-scale multimodal biomedical instruction-tuning dataset for mixed-modal foundation models. With 1.47 million instances, MedMax encompasses a diverse range of tasks, including multimodal content generation (interleaved image-text data), biomedical image captioning and generation, visual chatting, and report understanding. These tasks span diverse medical domains such as radiology and histopathology. Subsequently, we fine-tune a mixed-modal foundation model on the MedMax dataset, achieving significant performance improvements: a $26\%$ gain over the Chameleon model and an $18.3\%$ improvement over GPT-4o across 12 downstream biomedical visual question-answering tasks. Additionally, we introduce a unified evaluation suite for biomedical tasks, providing a robust framework to guide the development of next-generation mixed-modal biomedical AI assistants.
Code: https://github.com/Hritikbansal/medmax
Website: https://mint-medmax.github.io/
@@
figurec Refer to caption Figure 1: Average performance of multimodal models on twelve VQA tasks. Our MedMax instruction-tuned mixed-modal foundation model outperforms both open multimodal models (Chameleon, LLaVA-Med-v1.5, and Huatuo) and closed multimodal models (GPT-4o, GPT-4o-mini). This underscores the effectiveness of the MedMax dataset in training capable multimodal biomedical assistants.
00footnotetext: {*} Equal Contribution (alphabetical order).
@@ 1 Introduction
Recently, there has been rapid advancement in the development of mixed-modal foundation models that can perceive and generate data from multiple modalities such as Chameleon [54], Transfusion [76], Aurora [8], Gemini-2.0-Flash [18], and others [75, 58, 51, 70]. In particular, these models offer a flexible design, which embeds multiple modalities into a shared embedding space and uses a transformer backbone to learn interactions between these diverse modalities. During pretraining, these models are exposed to internet-scale data that equips them with the knowledge to perform real-world tasks involving multiple modalities within a unified architecture (e.g., image captioning and image generation). The ability of foundation models to understand and reason across the vast landscape of biomedical data—including multiple modalities (e.g., image-caption pairs), sources (e.g., medical research papers, YouTube channels), and domains (e.g., radiology, histology)—positions them as powerful tools for biomedical diagnosis, prognosis, and prevention [6, 59]. their multimodal generation capabilities can facilitate medical treatment planning by predicting a patient’s CT scan outcomes following specific treatments.
@@
However, existing mixed-modal foundation models struggle to perform well on vision-language biomedical data (e.g., visual question answering on X-rays) due to significant distribution shifts from the more commonly occurring natural data found on the internet (e.g., everyday objects and scenes). In this context, instruction tuning [60, 39, 29] offers a promising approach to understanding novel user intents and unlocking new capabilities for developing advanced biomedical assistants. But, there is a lack of large-scale multimodal biomedical instruction-tuning datasets, which are crucial for enabling mixed-modal models to reason and solve complex biomedical tasks across diverse domains.
@@
Traditional biomedical visual question answering (VQA) datasets such as VQA-RAD [30], SLAKE [36], and PathVQA [20] are crucial for imparting domain-specific knowledge. However, they suffer from several limitations such as lack of scale (e.g., typically only thousands of instances), task diversity (e.g., not suitable for teaching biomedical captioning or generation), and narrow focus on specific domains (e.g., radiology or pathology). In contrast, LLaVA-Med [39] collects biomedical vision-language alignment data (i.e., image-text pairs) and curated synthetic instruction-tuning datasets (i.e., image-conditioned conversational data). This aids in allowing the foundation models to assist the practitioners for novel queries about the biomedical images. However, it relies on PMC-15M [72], which is not openly available. Moreover, much of the LLaVA-Med data consists of figures and plots rather than biomedical images, impacting its overall quality.
Subsequent work like PubMedVision [71] aims to improve data quality through synthetic data curation across diverse biomedical domains. Nevertheless, it is limited to the knowledge available in medical research papers, restricting its ability to generate medical reports (e.g., findings and impressions based on biomedical imagery) that could assist doctors. Other datasets, such as Quilt-LLaVA, independently create visual instruction-tuning data for histopathology using alternative data sources, such as YouTube videos. Similarly, MIMIC-CXR provides expert-written radiology reports for chest X-rays. Furthermore, the ability to generate biomedical images is crucial for alleviate the lack of annotated medical imaging datasets [13] and disease progression modeling [19]. However, there are no existing biomedical instruction tuning datasets that focus on biomedical text-to-image generation and interleaved image-text content.
@@
Subsequently, we fine-tune a mixed-modal foundation model, Chameleon [54, 15], on the MedMax dataset. The dataset comprises a total of $1.7$B multimodal discrete tokens for instruction tuning. In our experiments, the MedMax-fine-tuned model outperforms the base Chameleon and GPT-4o [23] by $26\%$ and $18.3\%$ percentage points, respectively, when averaged across a set of 12 downstream VQA evaluation tasks (Figure 1). Furthermore, we conduct extensive experiments on diverse tasks that mixed-modal models excel in, including biomedical image captioning, image generation, visual chatting, and multimodal data generation. Given the general lack of support for biomedical evaluation, we provide a comprehensive evaluation suite encompassing diverse tasks to enable unified and efficient assessments. Overall, our work establishes a strong foundation for high-quality instruction tuning data creation, model fine-tuning, and robust evaluation of next-generation mixed-modal models.
Figure 2: Qualitative Examples. Sample generations from the MedMax model across diverse tasks: multimodal (interleaved) generation, visual chat, image captioning, and image generation.
2 MedMax
@@ 3.1 Mixed-Modal Foundation Models
By incorporating a diverse set of multimodal sequences at an internet scale, these pretrained models can achieve strong performance on downstream tasks such as image generation and image captioning. We illustrate the architecture for a mixed-modal foundation model in Figure 3.
@@ 3.2 Instruction Tuning
While internet-scale pretraining equips foundation models with diverse world knowledge, further instruction tuning is conducted to develop new skills and teach them how to interact with humans as assistants [60, 53]. In this context, it is essential to instruction-tune mixed-modal foundation models to impart task-specific skills. Formally, the instruction tuning data comprises paired multimodal sequences $(x,y)$, where $x$ represents the ‘instruction’ and $y$ represents the ‘response.’ Both $x$ and $y$ may include image tokens, text tokens, or a combination of both. For instance, in the context of visual question answering (VQA) tasks, $x$ includes an input image along with a corresponding question, while $y$ provides the correct response. The instruction tuning objective for a dataset $\mathcal{D}_{\mathcal{I}}$ can be formalized as:
@@
Unlike the pretraining objective, which models the entire sequence, the instruction tuning objective computes the loss solely on the ‘response’ portion of the sequence, thereby concentrating the learning process on the desired tasks.
@@ 4 MedMax
Due to the internet-scale pretraining and flexibility of the mixed-modal foundation models, we aim to imbibe the capability to understand human intents to solve diverse biomedical tasks (e.g., VQA, generation), domains (e.g., radiology, histopathology), and modalities (e.g., image and text content) through instruction tuning. Thus, we present MedMax, an instruction tuning data designed training mixed-modal foundation models for the biomedical applications. Specifically, the dataset construction involves: (a) designing a new instruction-tuning data that allows interleaved image-text content as a model response, (b) collecting various data sources to endow diverse skills into the model, and (c) performing dataset-specific curation to ensure high-quality. We illustrate the examples for diverse skills in Figure 4. We will go through each of these components in this section.
4.1 Biomedical Multimodal Generation Instruction Tuning (MedMax-Instruct)
With the emergence of mixed-modal foundation models, generating interleaved image-text content has become possible. Specifically, such multimodal output capabilities have several novel applications, including advanced diagnostics (e.g., visualizing the effects of specific treatments on biomedical image markers), generating patient reports that are often multimodal, and enhancing medical training by providing contextually relevant multimodal content for learners. However, no instruction-tuning datasets currently exist to equip mixed-modal assistants with this capability. To address this challenge, we present a multimodal output instruction-tuning data (MedMax-Instruct) for training biomedical mixed-modal assistants. Broadly, this dataset is created by prompting the LLM (GPT-4o) to generate a single-turn multimodal conversation based on the image description. Prior works such as PMC-VQA [73] and LLaVA [39] have utilized the capability of the LLMs to annotate high-quality biomedical data conditioned on the variables like image description. In particular, we create this dataset in three stages:
@@ 4.1 Biomedical Multimodal Generation Instruction Tuning (MedMax-Instruct)
Here, we intend to collect diverse biomedical image-text pairs that can be purposed for this task. In this regard, we combine image-caption instances from the PMC-OA and Quilt data. Specifically, our PMC-OA-specific curation (§4.2) reduces the original size of the PMC-OA (val split) from $166$K to $73$K instances. [^1] We reserve the use of PMC-OA (train split) for image captioning and generation task, as discussed in §4.2. In addition, we randomly choose $50$K instances from Quilt data which does not overlap with the Quilt examples in the prior subsets of the MedMax data. In total, we have $123$K image-text data which will be further filtered to ensure high-quality.
@@
Since we aim to generate multimodal conversations conditioned on the captions, it is critical to ensure that they describe the visual contents of the image well. To address this, we filter the captions that GPT-4o-mini finds to be of low quality. We present the prompt used for assessing the caption quality in Appendix Table 4. Following this, we are left with $88$K instances, leading to the removal of $25\%$ captions. We spent \$$\$12$ in performing this quality check using the API.
@@
Here, we prompt GPT-4o to generate single-turn multiodal conversation conditioned on the captions in the remaining dataset. We choose GPT-4o as it achieves state-of-the-art performance on text-only medical datasets such as MedQA [43]. We present our data generation template in Appendix Table 5. Specifically, we ask the GPT model to output an image placeholder (‘<image>’) which is then replaced with the ground-truth image from the image-caption pair.
In total, we spent \$$\$500$ to collect GPT-4o responses using the API. Finally, we have 88K instances consisting novel query (grounded in text) and multimodal response (interleaved text-image) for the biomedical applications. We emphasize that our data generation strategy is highly scalable and opt to limit the dataset size to current scale due to budget constraints. In addition, we focus on equipping the mixed-modal model with additional skills to enhance the diversity of the MedMax dataset.
Figure 5: Task-specific data sources in MedMax. We present the data-sources used to curate task-specific data in the MedMax collection. We abbreviate PubMedVision as PMV, visual question answering as VQA, Image captioning and generation tasks as Image Cap & Gen.
Figure 6: Distribution of skills (left) and data sources (right) in MedMax. We highlight that MedMax consists diverse skills for mixed-modal instruction tuning (left). In particular, we source the data from biomedical sources that cover several domains (e.g., radiology) and knowledge bases (e.g., research papers, YouTube).
@@ 4.2 Dataset Curation for Diverse Skills
We aim to collect a diverse set of biomedical datasets to teach novel skills to the mixed-modal foundation models. We curate high-quality datasets for individual skills as follows:
@@
The ability to answer questions for the images is crucial for biomedical assistants. In this regard, we include biomedical visual questions that involve answering close-ended questions (yes/no), open-ended questions, and multiple-choice questions in MedMax. In this context, the biomedical image and question constitute the input (or context), while the ground-truth answer serves as the output (or response).
@@
Prior work such as LLaVA-1.5 [38, 32] have included the data from the general-purpose VQA benchmarks in their instruction tuning mix to enhance the model’s visual capabilities. Hence, we combine the training sets of popular VQA datasets including VQA-RAD [30], SLAKE [36], PathVQA [20], and PMC-VQA [73]. To embed the knowledge about different anatomical regions ($20+$ regions), we split the OmnimedVQA [22] dataset into a training (81K) and testing set (1K), and add the training set into the MedMax data. In addition to the diversity in the domains and question styles, these datasets are also a rich source of expert-annotated data (e.g., clinician-driven annotation of VQA-RAD) that is critical for building high-quality data. In total, we have 284K VQA instances in the MedMax dataset.
Image captioning and generation: The ability to interpret and analyze the biomedical images is critical for accurate disease monitoring and diagnosis [46]. On the flip side, the capability to generate realistic biomedical images conditioned on the medical concepts, grounded in text, is useful for addressing the lack of high-quality annotated biomedical imaging data [13]. In this regard, a mixed-modal foundation model can benefit from potential knowledge transfer between both the tasks in the unified architecture. Here, we re-purpose a wide range of biomedical image-caption data found on the internet for the image captioning and generation task.
Table 1: Additional information about diverse biomedical dataset sources. We highlight that MedMax consists data across several biomedical domains and knowledge bases.
Data source Domain Knowledge Base
LLaVA-Med-PMC Diverse PubMed Central [2]
PMC-OA Diverse PubMed Central [2]
Quilt-1M Histopathology YouTube [3]
LLaVA-Med-IT Diverse PubMed Central [2]
PubMedVision-Alignment Diverse PubMed Central [2]
PubMedVision-IT Diverse PubMed Central [2]
Quilt-Instruct Histopathology YouTube [3]
VQA-RAD Radiology MedPix [1]
SLAKE Radiology MSD [7], CXR-8 [57], Chaos [28]
PathVQA Pathology PEIR Digital Library [27]
PMC-VQA Radiology PubMed Central [2]
OmniMedVQA Diverse Diverse
MIMIC-CXR Chest X-ray MIMIC-CXR [26]
@@
To this end, we collect image-caption data from four sources: (a) LLaVA-Med-PMC, (b) PMC-OA [35], (c) Quilt [24], and (d) PubMedVision-AlignmentVQA [71]. In particular, LLaVA-Med-PMC refers to the $600$K subset of PMC-15M [72] (curated from PubMed Central research papers) that was released as a part of LLaVA-Med data. On manual inspection, we found that this data is mostly composed of statistical figures, graphs and plots on biomedical data instead of biomedical images (e.g., X-ray, MRI, CT). Since the focus of our work is to train models on biomedical data, we filter this data using BioMedCLIP score (Appendix C). Subsequently, we were left with $37$K image-caption instances from LLaVA-Med-PMC data.
Further, we utilize the PMC-OA data that also consists of image-caption data curated from PubMedCentral research papers. Its notable features includes its accessibility (open-access), size ($1M+$ instances) and diversity of biomedical imaging data (e.g., ultrasound, fMRI, endoscope, PET). To maintain one-to-one correspondence in the data, we filter instances where a caption was aligned with multiple sub-images. Further, we filter small biomedical images that were less than 200 pixels in width or height in this data. Overall, the filtered training PMC-OA split contains $83$K instances.
While the previous datasets are constructed from the biomedical research papers, we also include Quilt [24] in our dataset. Specifically, it consists 1M image-text pairs of histopathology dataset that is largely curated from the lectures on the youtube. We utilize $100$K subset of this data for teaching for the captioning and generation subset of the MedMax data. Subsequently, we purpose these real image-text pairs for image captioning (image is the context and text is the response) and generation (text is the context and image is the response). We provide the finetuning templates in Appendix §D.
In many cases, the real image-caption has inherent data noise and formatting issues. As a result, we also include the synthetically-generated PubMedVision-AlignmentVQA data which utilizes GPT-4-Vision [44] to denoise and reformat internet data. In particular, we filter the original data ($647$K) to remove multiple image instances in this data to get $504$K instances, and randomly select a subset of $100$K instances for the MedMax mix. While this dataset can be utilized for image captioning (input image and question as context, and the answer as context), it could not be utilized for image generation directly. To address this, we prompt GPT-4o-mini [23] to convert the image descriptions in the PubMedVision-AlignmentVQA data into image generation prompts: ‘Convert the image description into an image generation prompt with AI’. We show an example in Appendix Table 6. In total, we curate 320K instances for the image captioning ($160$K) and the image generation ($160$K).
@@
Biomedical visual instruction tuning [39, 33] enables foundation models to reason and solve diverse tasks related to biomedical images for the practitioners. Hence, we aim to curate a wide range of tasks that practitioners care about for various biomedical domains (e.g., radiology, histopathology) and sources (e.g., research papers and internet videos). Specifically, we utilize the existing data sources such as: (a) LLaVA-Med-Instruct [33], (b) Quilt-Instruct [49], and (c) PubMedVision-Instruct [71]. In particular, these datasets are created synthetically by prompting foundation models such as LLMs [5] with image descriptions or external knowledge (e.g., inline mentions in research papers or video transcripts).
Concretely, we collect $76$K instances from LLaVA-Med-instruct-120K data. The remaining instances were not available publicly. In particular, half of this data was constructed using the image description and the other half also utilized the inline mentions of the images in the PubMedCentral research papers. To further enrich our data with diverse instructions, we include PubMedVision-IT dataset which is created using GPT-4-Vision [4]) as the data generator. Originally, this data consists $647$K instances, but we filter instances with multiple images in the context that left us with $504$K instances. Finally, we include $107$K conversations from the Quilt-Instruct [49] to diversify our dataset with knowledge from the youtube videos (and transcriptions) for the histopathology domain. Overall, the MedMax consists 686K instances for visual chat scenarios.
@@
The ability to perform detailed inspection of a patient’s imaging data requires specialized training. Thus, it is vital to expose our model to the expert-written findings (normal and abnormal anatomical cues in the image) from the patient’s data. Hence, we collect chest radiographs along with the medical reports in the MIMIC-CXR [26] data. Originally, the dataset consisted of 377K instances. We filtered it to exclude reports discussing more than one image, reducing the dataset to 102K instances. Subsequently, we subsampled the data to decrease the proportion of ‘No findings’ reports [13] from $20\%$ to $10\%$. Finally, we have 92K instances consisting chest radiograph-report pairs. We purpose half of the dataset for radiograph-conditioned report generation task, and the other half to generate chest radiographs conditioned on the medical report. We provide the templates used for these tasks in Appendix D (Table 10 and 11).
We present the details of the task-specific data sources in Figure 5. Further, we highlight the biomedical domains and knowledge bases covered in the MedMax in Table 1. Additionally, Figure 6 provides details on the biomedical domain, database, and the proportion of individual dataset sources. Furthermore, we note that MedMax comprises 725K unique images and 947K unique words. Collectively, these features underscore the diversity across various quality axes, making it well-suited for instruction tuning of mixed-modal foundation models.

### v2

@@ Abstract
Recent advancements in mixed-modal generative have opened new avenues for developing unified biomedical assistants capable of analyzing biomedical images, answering complex questions about them, and generating multimodal patient reports. However, existing datasets face challenges such as small sizes, limited coverage of biomedical tasks and domains, and a reliance on narrow sources. To address these gaps, we present MedMax, a large-scale multimodal biomedical instruction-tuning dataset for mixed-modal foundation models. With 1.47 million instances, MedMax encompasses a diverse range of tasks, including interleaved image-text generation, biomedical image captioning and generation, visual chat, and report understanding. These tasks span knowledge across diverse biomedical domains, including radiology and histopathology, grounded in medical papers and YouTube videos. Subsequently, we fine-tune a mixed-modal foundation model on the MedMax dataset, achieving significant performance improvements: a $26\%$ gain over the Chameleon model and an $18.3\%$ improvement over GPT-4o across 12 downstream biomedical visual question-answering tasks. Finally, we introduce a unified evaluation suite for biomedical tasks to guide the development of mixed-modal biomedical AI assistants. The data, model, and code is available at https://mint-medmax.github.io/.
00footnotetext: {*} Equal Contribution (alphabetical order).
@@
figurec
Figure 1: Average performance of multimodal models on twelve VQA tasks. Our MedMax instruction-tuned mixed-modal foundation model outperforms both open multimodal models (Chameleon, LLaVA-Med-v1.5, and Huatuo) and closed multimodal models (GPT-4o, GPT-4o-mini). This underscores the effectiveness of the MedMax dataset in training capable multimodal biomedical assistants.
@@ 1 Introduction
Recently, there has been rapid advancement in the development of mixed-modal foundation models that can perceive and generate data from multiple modalities such as Chameleon [34] and Transfusion [69]. During pretraining, these models are exposed to internet-scale data that equips them with the knowledge to perform real-world tasks involving multiple modalities within a unified architecture (e.g., image captioning and image generation). Their native multimodal capabilities have unlocked new opportunities to tackle challenging biomedical tasks, including analyzing patient scans for accurate diagnosis and generating multimodal medical reports [1, 53, 12].
@@
However, existing mixed-modal foundation models struggle to perform well on vision-language biomedical data due to significant distribution shifts from the more commonly occurring natural data found on the internet (e.g., everyday objects and scenes). In this context, instruction tuning [54, 30, 21] offers a promising approach to understanding novel user intents and unlocking new capabilities for developing advanced biomedical assistants. But, there is a lack of large-scale multimodal biomedical instruction-tuning datasets, which are crucial for enabling mixed-modal models to reason and solve complex biomedical tasks across diverse domains.
@@
Traditional biomedical visual question answering (VQA) datasets such as VQA-RAD [22], SLAKE [28], and PathVQA [13] are crucial for imparting domain-specific knowledge. However, they are limited in the dataset size (e.g., typically only thousands of instances). Other work like LLaVA-Med [30] collects biomedical vision-language alignment data (i.e., image-text pairs) and curated synthetic instruction-tuning datasets (i.e., image-conditioned conversational data). This aids in allowing the foundation models to assist the practitioners for novel queries about the biomedical images. However, much of the dataset consists of figures and plots rather than biomedical images, impacting its overall quality. Subsequently, PubMedVision [64] curated synthetic biomedical data curation using multimodal foundation models [38]. Nevertheless, the scope of their dataset is limited to the biomedical knowledge available in medical research papers while other works have shown significant biomedical knowledge is grounded in the youtube videos [45] as well as doctor-assisted medical reports [18]. In addition to conversational capabilities, a native multimodal model can aid in visualizing patient reports, enabling applications such as the creation of annotated medical datasets [7] and modeling disease progression [12]. However, no instruction-tuning datasets currently exist for the unified training of a multimodal biomedical assistant with diverse capabilities.
@@
Subsequently, we fine-tune a mixed-modal foundation model, Chameleon [34, 9], on the MedMax dataset. The dataset comprises a total of $1.7$B multimodal discrete tokens for instruction tuning. In our experiments, the MedMax-fine-tuned model outperforms the base Chameleon and GPT-4o [16] by $26\%$ and $18.3\%$ percentage points, respectively, when averaged across a set of 12 downstream VQA evaluation tasks (Figure 1). Given the general lack of support for biomedical evaluation, we provide a comprehensive evaluation suite encompassing diverse tasks to enable unified and efficient assessments. Thus, we conduct extensive experiments on diverse tasks that mixed-modal models excel in, including biomedical image captioning, image generation, visual chatting, and multimodal generation. Overall, our work establishes a strong foundation for high-quality instruction tuning data creation, model fine-tuning, and robust evaluation of next-generation mixed-modal models.
@@ 2.1 Mixed-Modal Foundation Models
By incorporating a diverse set of multimodal sequences at an internet scale, these pretrained models can achieve strong performance on downstream tasks such as image generation and image captioning (Figure 3).
@@ 2.2 Instruction Tuning
While internet-scale pretraining equips foundation models with diverse world knowledge, further instruction tuning is conducted to develop new skills and teach them how to interact with humans as assistants [54, 48]. In our context, the instruction tuning data comprises paired multimodal sequences $(x,y)$, where $x$ represents the ‘instruction’ and $y$ represents the ‘response.’ Both $x$ and $y$ may include image tokens, text tokens, or a combination of both. For instance, in the context of visual question answering (VQA) tasks, $x$ includes an input image along with a corresponding question, while $y$ provides the correct response. The instruction tuning objective for a dataset $\mathcal{D}_{\mathcal{I}}$ can be formalized as:
@@
In particular, the instruction tuning objective computes the loss solely on the ‘response’ portion of the sequence, thereby concentrating the learning process on the desired tasks.
@@ 3 MedMax
In this work, we aim to solve diverse biomedical tasks (e.g., VQA, generation), domains (e.g., radiology, histopathology), and modalities (e.g., image and text content) through instruction tuning. Thus, we present MedMax, an instruction tuning data designed training mixed-modal foundation models for the biomedical applications. Specifically, the dataset construction involves: (a) designing a new instruction-tuning data that allows interleaved image-text content as a model response (§3.1), and curating various data sources to endow diverse skills into the model (§3.2).
Figure 4: We present the data-sources used to curate task-specific data in the MedMax collection.
Figure 5: We source the data from biomedical sources that cover several domains (e.g., radiology) and knowledge bases (e.g., research papers, YouTube).
3.1 MedMax-Instruct
With the rise of mixed-modal foundation models, generating interleaved image-text content is now possible. This capability enables novel applications like advanced diagnostics (e.g., visualizing treatment effects on biomedical markers), generating multimodal patient reports, and enhancing medical training with relevant multimodal content. However, no instruction-tuning datasets exist for training mixed-modal assistants in this area. Therefore, we introduce MedMax-Instruct, a multimodal generation instruction-tuning dataset for biomedical mixed-modal assistants, created in three stages as described:
@@ 3.1 MedMax-Instruct
Here, we intend to collect diverse biomedical image-text pairs that can be purposed for this task. In this regard, we combine image-caption instances from the PMC-OA and Quilt data. Specifically, our PMC-OA-specific curation (§3.2) reduces the original size of the PMC-OA (val split) from $166$K to $73$K instances. [^1] We reserve the use of PMC-OA (train split) for image captioning and generation task, as discussed in §3.2. In addition, we randomly choose $50$K instances from Quilt data. In total, we have $123$K image-text data which will be further filtered to ensure high-quality.
@@
Since we aim to generate multimodal conversations conditioned on the captions, it is critical to ensure that they describe the visual contents of the image well. To address this, we filter the captions that GPT-4o-mini [39] finds to be of low quality. We present the prompt used for assessing the caption quality in Appendix Table 4. Following this, we are left with $88$K instances, leading to the removal of $25\%$ captions.
@@
Here, we prompt GPT-4o to generate single-turn multiodal conversation conditioned on the captions in the remaining dataset. We choose GPT-4o as it achieves state-of-the-art performance on text-only medical datasets such as MedQA [36]. We present our data generation template in Appendix Table 5. Specifically, we ask the GPT model to output an image placeholder (‘<image>’) which is then replaced with the ground-truth image from the image-caption pair. In total, we spent \$$\$500$ to collect GPT-4o responses using the API. Finally, we have 88K instances consisting novel query (grounded in text) and multimodal response (interleaved text-image) for the biomedical applications. We present an example for multimodal generation in Figure 2. Our data generation strategy is highly scalable, and we have limited the dataset size due to budget constraints.
@@ 3.2 Dataset Curation for Diverse Skills
Further, we aim to specialized biomedical datasets to teach novel skills to the mixed-modal foundation models:
@@
We include biomedical visual questions that involve answering close-ended questions (yes/no), open-ended questions, and multiple-choice questions in MedMax. Prior work such as LLaVA-1.5 [29, 24] have included the data from the general-purpose VQA benchmarks in their instruction tuning mix to enhance the model’s visual capabilities. Hence, we combine the training sets of popular VQA datasets including VQA-RAD [22], SLAKE [28], PathVQA [13], and PMC-VQA [66]. To embed the knowledge about different anatomical regions ($20+$ regions), we split the OmnimedVQA [15] dataset into a training (81K) and testing set (1K), and add the training set into the MedMax data. In addition to the diversity in the domains and question styles, these datasets are also a rich source of expert-annotated data (e.g., clinician-driven annotations). In total, we have 284K VQA instances in MedMax dataset.
@@
Image captioning and generation: The ability to interpret and analyze biomedical images is crucial for accurate disease monitoring and diagnosis [41]. Additionally, the ability to generate images is valuable for creating high-quality annotated biomedical imaging data [7]. To this end, we first curate LLaVA-Med-PMC that refers to the $600$K subset of PMC-15M [65] that was released as a part of LLaVA-Med data. On manual inspection, we found that this data is mostly composed of statistical figures, graphs and plots instead of biomedical reports (e.g., X-ray, MRI, CT). To filter this data, we utilize BioMedCLIPScore [65] score (Appendix A.1). Subsequently, we are left with $37$K image-caption instances from LLaVA-Med-PMC.
Further, we utilize the PMC-OA [27] data that also consists of image-caption data curated from PubMedCentral research papers. We provide more details about its curation in Appendix A.2. Overall, the filtered training PMC-OA split contains $83$K instances. Further, we include $100$K subset of Quilt [17] histopathology dataset collected from youtube for the biomedical image captioning and generation tasks. Finally, we also include the synthetically-generated PubMedVision-AlignmentVQA [64] data which utilizes GPT-4-Vision [38] to denoise and reformat noisy internet data for biomedicine. We provide more details about its curation in Appendix A.3. In total, we curate 320K instances for the image captioning ($160$K) and the image generation ($160$K).
@@
Therefore, we aim to curate a diverse set of queries related to biomedical images that are relevant to practitioners across various biomedical domains (e.g., radiology, histopathology) and sources (e.g., research papers, internet videos). Concretely, we collect $76$K instances from LLaVA-Med-instruct-120K [25] data. The remaining instances were not available publicly. In particular, half of this data was constructed using the image description and the other half also utilized the inline mentions of the images in the PubMedCentral research papers. To further enrich our data with diverse instructions, we include PubMedVision-IT [64] dataset which is created using GPT-4-Vision [37]). Originally, this data consists $647$K instances, but we filter instances with multiple images in the context that left us with $504$K instances. Finally, we include $107$K conversations from the Quilt-Instruct [45] to diversify our dataset with knowledge from the youtube videos (and transcriptions) for the histopathology domain. Overall, the MedMax consists 686K instances for visual chat scenarios.
@@
The ability to perform detailed inspection of a patient’s imaging data requires specialized training. Thus, it is vital to expose our model to the expert-written findings (normal and abnormal anatomical cues in the image) from the patient’s data. Hence, we collect chest radiographs along with the medical reports in the MIMIC-CXR [18] data. We provide the details about its curation in Appendix A.4. Finally, we have 92K instances consisting chest radiograph-report pairs. We purpose half of the dataset for radiograph-conditioned report generation task, and the other half to generate chest radiographs conditioned on the medical report. We provide the templates in Appendix D (Table 10 and 11).
We illustrate examples for diverse skills included in the MedMax in Figure 2. In addition, we illustrate the details of the task-specific data sources in Figure 4. Further, we highlight the biomedical domains and knowledge bases covered in the MedMax in Table 3. Additionally, Figure 5 provides details on the biomedical domain, database, and the proportion of individual dataset sources. Overall, we note that MedMax comprises 725K unique images and 947K unique words. Collectively, these features underscore the diversity across various quality axes, making it well-suited for instruction tuning of mixed-modal foundation models.